how long is each side of a square if the diagonal is 5√2

Answer by stanbon(75887)

how long is each side of a square if the diagonal is 5√2
------
Let each side be "x":
x^2 + x^2 = (5sqrt(2))^2
-------
2x^2 = 25*2
x^2 = 25
----
side = x = 5
=================
Cheers,
Stan H.
=================